A rectangle with length of 8 units and a width 6 units. whats is the length of a diagonal
Answer by mananth(16946)

The sides & the diagonal from a right triangle. diagonal = hypotenuse.
legs are 6 & 8
..
6^2+8^2= D^2
36+64=D^2
100=D^2
take the square root
D= 10
...
m.ananth@hotmail.ca